OpenStudy (anonymous):

PLEASE HELP. -y^2+4x-10y-1=0 Write in standard form.

OpenStudy (anonymous):

what do you need help with writing it in standrard form?

OpenStudy (anonymous):

the equation, it says classify each conic section and write in standard form by completing the square. i got the square i just need help writing it out in standard form.

OpenStudy (anonymous):

Ask @perl im sure he will help you with this

OpenStudy (anonymous):

Ax By=c

OpenStudy (amoodarya):

\[(y-b)^2=4p(x-a)\] it is standard form

OpenStudy (amoodarya):

\[-y^2+4x-10y-1=0\\4x=y^2+10y+1\\4x=(y+5)^2-5^2+1\] can you go on ?

OpenStudy (anonymous):

okay so thats the answer right 4x=(y+5)^2-5^2+1

OpenStudy (anonymous):

?

OpenStudy (amoodarya):

\[(y+5)^2-25+1=4x\\(y+5)^2=4x+24\\(y+5)^2=4(x+6)\]

OpenStudy (amoodarya):

that's it